Calculate the acceleration of Josh riding his bicycle in a straight line that speeds up from 4m/s to 6m/s in 5 seconds.


1
Expert's answer
2022-05-16T13:41:30-0400
"a=\frac{\Delta v}{\Delta t}=\rm\frac{6\: m/s-4\: m/s}{5\: s}=0.4\: m/s^2"